# Training Small Reasoning LLMs with Cognitive Preference Alignment

### v1

@@ 4.4 Comparison Against Other Models
We compare our trained 7B model with other models released in the open-source community. The DeepSeek team DeepSeek-AI (2025) releases distilled models based on a (significantly larger) 800K proprietary SFT dataset, which is not publicly available. To ensure fair comparisons, we consider two reasoning LLMs available before the launch of DeepSeek-R1, namely Macro-o1 Zhao et al. (2024) and LLaMA-O1131313https://github.com/SimpleBerry/LLaMA-O1. We also compare other models trained on datasets distilled from DeepSeek-R1 in the open-source community, including Bespoke-Stratos-7B [^14] https://huggingface.co/bespokelabs/Bespoke-Stratos-7B and OpenThinker-7B [^15] https://huggingface.co/open-thoughts/OpenThinker-7B. The results, along with the sizes of training sets, are shown in Table 2. The results indicate that models trained on datasets distilled from DeepSeek-R1 perform significantly better compared to others.
Using our approach, we train two models using the Bespoke-Stratos-17k and OpenThoughts-114k training sets, respectively. Thus, it is fair to compare our method against that of Bespoke-Stratos-7B and OpenThinker-7B. As seen, our resulting models demonstrate significantly better performance compared to baselines in all benchmarks, which clearly proves the effectiveness of our approach. By leveraging larger training sets, we can surely obtain stronger models based on our approach, which will be left as future work.
@@ 4.5 Study on Choices of the Critic
In the previous section, we have claimed that using the small target LLM $\pi_{\text{base}}$ as the Critic does not necessarily produce satisfactory results due to the limited parameter size. In contrast, larger LLMs $\pi_{\text{large}}$ can better “think” like small models. The results of CRV using different backbones as the Critic are shown in Table 3, with the backbones for the Rethinker and the Verifier unchanged. From the results, we can see that they confirm our findings, where larger models consistently perform better than the 7B model in most tasks. Among the three large agents, DeepSeek-R1-Distill-Qwen-32B exhibits the best performance based on majority voting across all the testing sets. Note that we do not leverage the QwQ model series as the Rethinker (instead we use the original Qwen instruct models) because it sometimes generates recursive reasoning loops, not suitable to serve our purpose.
@@ 4.6 Study on Model Scales
To study the effectiveness of different parameter sizes on the student models, we further report the performance of Qwen2.5-3B-Instruct and Qwen2.5-14B-Instruct. The experimental settings are identical to those of Qwen2.5-7B-Instruct. The results are presented in Figure 4. We observe that our method is also effective across different model scales. An interesting observation is that the improvement is more significant in Qwen2.5-14B-Instruct compared to Qwen2.5-3B-Instruct. This is because even when we leverage the CRV system to rewrite the CoTs, the large capacity gap between the teacher and student models makes it more challenging for Qwen2.5-3B-Instruct to capture the CoTs by SFT. This finding is also consistent with the recently discovered “distillation scaling law” Busbridge et al. (2025).
@@
Figure 4: Experimental results of different sizes of Qwen2.5 models on AIME2024, MATH500, GPQA Diamond, and LiveCodeBench V2.
@@ 4.7 Study on Other Model Backbones
To evaluate the universality of the proposed approach, we perform additional experiments on multiple backbone models beyond the Qwen2.5 series on Bespoke-Stratos-17k dataset. Figure 5 demonstrates that, for both LLaMA and Mistral models, our approach achieves notable performance gains over the direct SFT baseline across diverse mathematical and coding tasks. These results indicate that the CRV+CogPO framework enables seamless adaptation to different LLM backbones, demonstrating the universality of our approach on various LLM backbones, which also shows the potential of our work to produce stronger models based on other LLMs.
@@
Figure 5: Experimental results of different model series (Llama3.1-8B-Instruct, Mistral-7B-V0.3) other than Qwen2.5 on AIME2024, MATH500, GPQA Diamond, and LiveCodeBench V2.
Figure 6: Impact of different $\beta$ on AIME2024, GPQA Diamond, LiveCodeBench V2 and OlympiadBench.
@@ 4.8 Hyper-parameter Analysis
To evaluate the impact of $\beta$ values in CogPO, we perform a series of experiments with varying $\beta$ values to assess the algorithm’s effectiveness. As shown in Figure 6, the highest performance is attained when assigning tailored $\beta$ values to samples based on their specific gaps, which is a core principle of the CogPO algorithm.
@@ 4.9 Case Studies
Due to space limitations, case studies of our model outputs are shown in the appendix. They clearly show how our approach can effectively expand or simplify the reasoning processes based on the critic judgment for the training of small models.
@@ 5 Conclusion and Future Work
In this paper, we present the CRV framework, a multi-agent system that effectively trains smaller reasoning LLMs. Our approach leverages the strengths of specialized LLM agents to critique, refine, and verify CoT outputs, leading to enhanced reasoning capabilities in smaller models. The CogPO algorithm further aligns model outputs with their inherent cognitive capacities, significantly improving the performance on several challenging reasoning tasks. In the future, we will i) train and release stronger small models using larger CoT datasets; ii) improve the effectiveness of the CRV framework, especially for much smaller models; and iii) investigate our approach for other domain-specific applications, such as medical diagnosis and legal reasoning.
@@
While our proposed framework shows promising results in enhancing the reasoning capabilities of smaller LLMs, several limitations still remain. The CRV framework relies heavily on the contributions of larger models refining the CoT output. This dependency may create challenges in situations where access to larger models is restricted, or these larger models generate incorrect results. In addition, although our framework is designed for smaller LLMs, there remains a ceiling on their performance. By nature, smaller models inherently have reduced capacity to encode complex information and handle nuanced reasoning tasks, which may limit their effectiveness in certain scenarios.
@@
Our work is fully methodological; hence, there are no direct ethical issues. However, smaller models trained on data distilled from larger ones might inherit or exacerbate biased outputs, which can still influence outcomes. We suggest that continuous evaluation of trained LLMs based on ethical guidelines is indispensable.
Appendix A Supplementary Experiments and Discussion
A.1 Results on Weaker CoT Dataset
To demonstrate that our approach is truly superior to vanilla SFT over CoT datasets, we conduct an experiment on the Sky-T1 dataset, which is relatively weaker than Bespoke-Stratos-17k due to the choice of the teacher model (i.e., QwQ-32B-Preview) and the data curation pipeline. The results are presented in Table 4. As shown, in some cases, the SFT baseline cannot even beat the zero-shot performance because the 7B backbone is already strong. Hence, training on Sky-T1 with our small LLM backbone may not improve reasoning abilities. This negative finding is also consistent with their own blog regarding the model size and data quality [^16] https://novasky-ai.github.io/posts/sky-t1/. Nonetheless, by comparing our method with the SFT baseline, we can observe clear improvement, which demonstrates that our approach has the efficacy to enhance the reasoning abilities of small models in various scenarios.
Dataset/Model Zero-shot SFT Ours
AIME2024 10.0 16.7 20.0
MATH-500 73.6 73.2 77.0
GPQA Diamond 33.3 28.8 36.9
LiveCodeBench V2 30.7 20.9 23.9
A.2 Results on Larger CoT Dataset
Dataset/Model Zero-shot SFT Ours
AIME2024 10.0 31.3 43.3
MATH-500 73.6 83.0 88.4
GPQA Diamond 33.3 42.4 42.9
LiveCodeBench V2 30.7 39.9 46.4
A.3 Design Choice of the Critic
An initial, straightforward method is to employ $\pi_{\text{base}}$ as the Critic. Unfortunately, due to the limited reasoning capabilities inherent to the small model, it consistently faces difficulties in distinguishing the complexity levels effectively in our datasets. Consequently, we utilize a more robust and larger LLM for both the Rethinker and Verifier (referred to as $\pi_{\text{large}}$), which then functions as the Critic. This involves guiding the large model to simulate the cognitive approach of the smaller model, $\pi_{\text{base}}$. The prompt template of the Critic is shown in Table 10. This setting is akin to educational practices, where professors, instead of students, customarily curate academic content across a spectrum of difficulty levels due to their broader knowledge base. As shown in Table 3, the experiments clearly demonstrate the superior evaluative proficiency of the large model, confirming its advantage in categorizing CoT complexity from the perspective of the smaller model efficiently.
Another design choice is that, we divide the CoT processes into three categories: “easy”, “medium” and “hard”, instead of two categories: “correct” and “incorrect”. This is because for incorrect CoT processes, this binary classification fails to distinguish the root cause of errors (i.e., whether the CoT is overly simplified or overly complex), which is critical for determining appropriate refinement strategies (e.g., expansion for overly simplified processes vs. simplification for overly complex ones). Determining the correctness of CoT processes should be the responsibility of the Verifier.
Appendix B Testing Set Statistics
The sizes of our testing sets are summarized in Table 6.
@@ Appendix B Testing Set Statistics
Appendix C Hyperparameter Settings and Training Time
The default hyperparameters for model training and baselines are shown in Table 7. On the Bespoke-Stratos-17k dataset, for the 3B model, we use a single node with 8 A800 GPUs (80GB), with an approximate training time of 4 hours. For the 7B model, we use a single node with 8 A800 GPUs (80GB), with a training time of about 5 hours. For the 14B model, we use 4 nodes, each with 8 A800 GPUs (80G), resulting in a training time of approximately 14 hours.
@@ Appendix C Hyperparameter Settings and Training Time
Incorrect reasoning process Area=12(Sum of $\text{Area}=\frac{1}{2}(\text{Sum of sides})=\frac{1}{2}(3+4+5)=6$
@@
Your task is to assess the given reasoning process from the perspective of a small language model (e.g., 7B).
Specifically, determine whether the reasoning process provides sufficient detail for a small model to solve the
problem, or whether it is too simplistic (i.e., lacking critical details) or too complex (i.e., containing
@@
Difficulty Definitions (from the perspective of a small model):
- Easy: The reasoning process is overly simplistic relative to the problem’s difficulty; it omits essential details that
a small model needs to solve the problem.
@@
- Hard: The reasoning process is overly complex, with extraneous or convoluted steps that could hinder a small
model’s ability to follow it.
@@
You must output exactly one word: easy, medium, or hard. Do NOT provide any additional text, explanation.
@@
(easy) Given a problem ,its answer and its reasoning process, your task is to extend the reasoning process by adding
necessary details and intermediate steps so that a small language model (e.g., a 7B model) can follow the
extended reasoning process to solve the problem.
If the original reasoning process is divided into multiple steps separated by two newline characters, your output
must preserve this formatting.
@@
(hard) Given a problem, its answer and its reasoning process, your task is to simplify the reasoning process so that
a small language model (e.g., a 7B model) can reliably follow the steps to solve the problem.
If the original reasoning process is divided into multiple steps separated by two newline characters, your output
must preserve this formatting.
@@
Your task is to assess a given reasoning process based on a problem, its answer, and the reasoning process.
Specifically, determine whether the reasoning process is sufficient and accurate for a small model (e.g., 7B)
to reach the correct answer.
If the reasoning process aligns with the model’s cognitive abilities and can correctly guide the small model
to the solution, output YES.
If the reasoning process is flawed or does not match the model’s capabilities such that it fails to guide
the model to the correct answer, output NO.
You must output exactly one word: YES or NO. Do NOT provide any additional text, explanation.
Table 10: Prompt templates for the CRV system.
Appendix D Case Studies
Case studies are presented in Table 8 and Table 9.
Appendix E Prompt Templates
Prompt templates of the Critic, the Rethinker and the Verifier in our CRV system are shown in Table 10.

### v2

@@ 4.2 Comparison Against Other Models
We compare our trained 7B model with other models released in the open-source community. We consider two reasoning LLMs available before the launch of DeepSeek-R1, namely Macro-o1 Zhao et al. (2024) and LLaMA-O1888https://huggingface.co/SimpleBerry/LLaMA-O1-Supervised-1129. We also compare with models trained on datasets distilled from DeepSeek-R1, including Bespoke-Stratos-7B [^9] https://huggingface.co/bespokelabs/Bespoke-Stratos-7B and OpenThinker-7B [^10] https://huggingface.co/open-thoughts/OpenThinker-7B. Using our CRV+CogPO framework, we additionally train two models on the Bespoke-Stratos-17k and OpenThoughts-114k training sets, respectively. Thus, it is fair to compare our method against Bespoke-Stratos-7B and OpenThinker-7B. The results, along with the sizes of the training sets, are shown in Table 3. It can be observed that employing DeepSeek-R1-generated CoT data yields superior results. At the algorithmic level, both Bespoke-Stratos-7B and our model are trained on the 17K CoTs from DeepSeek-R1. Under identical data conditions, our model significantly outperforms Bespoke-Stratos-7B across all benchmarks and achieves performance comparable to OpenThinker-7B, which is trained on 114K CoTs from DeepSeek-R1. Moreover, when trained on the same dataset as OpenThinker-7B, our model substantially surpasses OpenThinker-7B on all benchmarks. These findings demonstrate that, given the same data, our CRV+CogPO training framework exhibits superior performance, confirming its effectiveness.
Dataset/Model Easy Medium Hard
AIME2024 13.3 23.3 16.7
MATH500 75.4 82.8 78.2
GPQA-D 34.3 37.4 33.3
LCB V2 31.9 36.2 32.5
Table 5: Experimental results on training data of different complexity levels.
@@ 4.3 Study on Choices of the Critic
In the previous section, we claimed that using the small target LLM $\pi_{\text{base}}$ as the Critic does not necessarily produce satisfactory results due to its limited parameter size. In contrast, larger LLMs $\pi_{\text{large}}$ can “think like small models” better. The results of using different backbones as the Critic are shown in Table 4, with the backbones for the Rethinker and the Verifier unchanged. From the results, we see that they confirm our findings, as larger models consistently perform better than the 7B model in almost all tasks. Among the three large agents, DeepSeek-R1-Distill-Qwen-32B exhibits the best performance based on majority voting across all testing sets. A detailed and in-depth analysis of the selection of the Critic is provided in Appendix A.5.
4.4 Training with CoT Datasets of Different Complexity Levels
To further investigate whether medium-level data are indeed the most suitable for the base model, we conduct experiments on the OpenThoughts-114K dataset. We use the Critic to rate all CoTs in the dataset, then randomly sampled 10K CoTs from each of the derived easy, medium, and hard subsets to construct three training sets. We then perform SFT with Qwen2.5-7B-Instruct on these three training sets under identical configurations. The results are shown in Table 5, indicating that when the number of training data is the same, the model trained on the medium subset achieves the highest scores, fully supporting our hypothesis. The CoTs in the easy and hard sets are either too terse or overly complex, preventing the base model from effectively comprehending all CoTs in those sets. In contrast, the medium-level subset aligns with the model’s cognitive capabilities and thus yields the best results.
Dataset/Model SFT w. C w. CR w. CRV
AIME2024 20.0 23.3 26.7 26.7
MATH500 80.0 83.4 83.2 84.0
GPQA-D 37.4 38.4 39.9 40.9
LCB V2 31.3 34.3 34.1 34.4
Table 6: Ablation results on the CRV system.
4.5 Study on Effectiveness of Critic, Rethinker, and Verifier
To further explore the collaborative mechanism within the CRV system and the individual roles and contributions of each component, we conduct extensive ablation experiments on the Bespoke-Stratos-17k dataset. Table 6 presents our ablation results. The “SFT” row reports results from directly performing SFT on the original dataset without any CRV intervention; the “w. C” row shows performance when only the Critic is applied before SFT, using only the traces rated as medium by the Critic for SFT; the “w. CR” row indicates results when both the Critic and the Rethinker participate prior to SFT, utilizing the medium-rated traces and the refined easy/hard traces that have not yet been verified; the “w. CRV” row reflects outcomes when the Critic, Rethinker, and Verifier are all applied.
As the Critic, Rethinker, and Verifier are added sequentially, the model’s reasoning ability exhibits a progressively improving trend, which clearly illustrates the role of each component. Notably, “w. CR” experiences a performance drop on MATH500 and LCB V2, indicating that omitting the Verifier after the Rethinker’s refinement can impair the model’s reasoning ability. Therefore, each component of the CRV system plays an indispensable role. To achieve optimal performance, we recommend processing the data using the complete system.
@@ 4.6 Study on Model Scales
To study the effectiveness of different parameter sizes on student models, we further report the performance of Qwen2.5-3B-Instruct and Qwen2.5-14B-Instruct. The experimental settings are identical to those of Qwen2.5-7B-Instruct. The results are presented in Figure 4. We observe that our method is also effective across different model scales. An interesting observation is that the improvement is more significant for Qwen2.5-14B-Instruct compared to Qwen2.5-3B-Instruct. This is because, even when we leverage the CRV system to rewrite the CoTs, the large capacity gap between the teacher and student models makes it more challenging for Qwen2.5-3B-Instruct to capture the CoTs through SFT. This finding is also consistent with the recently discovered “distillation scaling law” Busbridge et al. (2025).
@@
Figure 4: Experimental results for different sizes of Qwen2.5 models on AIME2024, MATH500, GPQA Diamond, and LiveCodeBench V2.
@@ 4.7 Study on Other Model Backbones
To evaluate the generality of the proposed approach, we perform additional experiments on multiple backbones beyond the Qwen2.5 series using the Bespoke-Stratos-17k dataset. Figure 5 demonstrates that, for both the LLaMA and Mistral series, our approach achieves notable performance gains over the direct SFT baseline across diverse mathematical and coding tasks. These results indicate that the CRV+CogPO framework enables seamless adaptation to other backbones, demonstrating the universality of our approach and its potential to produce stronger models based on other LLMs.
@@
Figure 5: Experimental results for other model series (Llama3.1-8B-Instruct, Mistral-7B-V0.3) beyond Qwen2.5, on AIME2024, MATH500, GPQA Diamond, and LiveCodeBench V2.
Figure 6: The study on impact of different $\beta$ values on AIME2024, GPQA Diamond, LiveCodeBench V2, and OlympiadBench.
4.8 Study on Computation Cost
Our CRV+CogPO framework consists of two main stages: data processing (the CRV system) and model training (SFT and CogPO). For model training, a detailed breakdown of the computational cost is provided in Appendix A.2. For data processing, the computational cost is primarily determined by use of the CRV system, which we explore in two key sections. Section 4.3 shows that models with stronger reasoning abilities yield superior results, while Section 4.5 confirms that every component of the CRV system positively contributes to the model’s reasoning performance.
Therefore, when computational resources are ample, we recommend selecting a powerful reasoning model as the agent and utilizing the complete CRV system to construct a high-quality training dataset for optimal performance. However, if computational resources are limited, our framework can still be effectively used to enhance reasoning capabilities. The experiments in Sections 4.3 and 4.5 show that even when employing a weaker model as the Critic or using only a partial CRV system, the resulting model’s reasoning ability still outperforms the original baseline.
@@ 4.9 Hyper-parameter Analysis
To evaluate the impact of $\beta$ values in CogPO, we perform a series of experiments with varying $\beta$ values to assess the algorithm’s effectiveness. As shown in Figure 6, the highest performance is attained when assigning tailored $\beta$ values to samples based on their respective gaps, which is a core principle of the CogPO algorithm.
@@ 4.10 Case Studies
Due to space limitations, case studies are included in the Appendix. These studies clearly show how our approach can effectively expand or simplify the reasoning processes based on the Critic’s feedback.
@@ 5 Conclusion and Future Work
In this paper, we present the CRV framework where we leverage the strengths of LLM agents to critique, refine, and verify CoT outputs for optimizing CoT training sets. The CogPO algorithm further aligns model outputs with their inherent cognitive capacities, improving performance on several challenging reasoning tasks. In the future, we will (i) train and release stronger small models using larger CoT datasets, (ii) improve the effectiveness of the CRV framework, especially for much smaller models, and (iii) investigate our approach for other domain-specific applications, such as medical diagnosis and legal reasoning.
@@
While our proposed framework shows promising results in enhancing the reasoning capabilities of smaller LLMs, several limitations remain. The CRV framework relies heavily on the contributions of larger models in refining the CoT outputs. This dependency may create challenges in situations where access to larger models is restricted, or when these larger models generate incorrect results. In addition, although our framework is designed for smaller LLMs, there remains a ceiling on their performance. By nature, smaller models inherently have reduced capacity to encode complex information and handle nuanced reasoning tasks, which may limit their effectiveness in certain scenarios.
@@
Our work is fully methodological; hence, there are no direct ethical issues. However, smaller models trained on data distilled from larger ones might inherit or exacerbate biased outputs, which can influence outcomes. We suggest that continuous evaluation of trained LLMs based on ethical guidelines is indispensable.
Acknowledgments
This work was supported by Alibaba Research Intern Program.
Appendix A Supplementary Experiments
A.1 Datasets
In our experiments, we evaluate our work on several benchmarks, including AIME2024111111https://huggingface.co/datasets/Maxwell-Jia/AIME_2024, MATH-500 Lightman et al. (2023), GSM8K Cobbe et al. (2021), GPQA Diamond Rein et al. (2023), LiveCodeBench V2 Jain et al. (2024), MMLU Hendrycks et al. (2021), and OlympiadBench (math-en) He et al. (2024). The sizes of our testing sets are summarized in Table 7.
For our training set $\mathcal{D}_{\text{SFT*}}$, we leverage Bespoke-Stratos-17k [^12] https://huggingface.co/datasets/bespokelabs/Bespoke-Stratos-17k, which contains 17K tuples of questions, reasoning processes, and answers directly distilled from DeepSeek-R1 DeepSeek-AI (2025). We also utilize two released CoT datasets to conduct supplementary experiments. The first is Sky-T1-data-17k [^13] https://github.com/NovaSky-AI/SkyThought, which is distilled from QwQ-32B-Preview—its reasoning abilities are reported to be weaker than those of DeepSeek-R1. The second is OpenThoughts-114k [^14] https://huggingface.co/datasets/open-thoughts/OpenThoughts-114k, which is distilled from DeepSeek-R1 and verified using a data curation recipe. We have chosen not to use some previously released CoT datasets (e.g., OpenLongCoT-SFT [^15] https://huggingface.co/datasets/SimpleBerry/OpenLongCoT-SFT) due to their significantly weaker reasoning abilities, while some benchmarks (e.g., AIME2024, OlympiadBench) are extremely challenging.
@@ A.1 Datasets
A.2 Experimental Details
In our work, we utilize Qwen2.5-7B-Instruct as the default model backbone and extend our evaluation to Llama3.1-8B-Instruct Dubey et al. (2024) and Mistral-7B-Instruct-v0.3 Jiang et al. (2023), along with other sizes of Qwen2.5 models, to validate the generalizability of our algorithm across diverse model architectures and sizes. We first establish a baseline by assessing the model’s zero-shot capabilities. Subsequent experiments leverage this result to quantify the performance improvements attributable to CRV and CogPO. During the CRV phase, the same generation hyperparameters are applied to the Critic, Rethinker, and Verifier for inference: temperature $T=0.7$, $\text{top\_p}=0.9$, and $\text{top\_k}=50$. The default backbone is DeepSeek-R1-Distill-Qwen-32B, while we test other backbone choices in the experiments. For CogPO training, the default $\beta$ values are: $\beta_{\text{S}}=0.1$, $\beta_{\text{M}}=0.2$, and $\beta_{\text{L}}=0.5$. Training details for all models and baselines are shown in Table 8.
On the Bespoke-Stratos-17k dataset: for the 3B model we use a single node with 8 A800 GPUs (80GB), with a training time of approximately 4 hours. For the 7B model, we use a single node with 8 A800 GPUs (80GB), with a training time of about 5 hours. For the 14B model, we use 4 nodes (each with 8 A800 GPUs), resulting in a training time of approximately 14 hours.
@@ A.2 Experimental Details
A.3 Results on Weaker CoT Dataset
To demonstrate that our approach is truly superior to vanilla SFT over CoT datasets, we conduct an experiment on the Sky-T1 dataset, which is relatively weaker than Bespoke-Stratos-17k due to the choice of teacher model (i.e., QwQ-32B-Preview) and the data curation pipeline. The results are presented in Table 9. As shown, in some cases the SFT baseline cannot even beat the zero-shot performance. This observation is consistent with their blog regarding model size and data quality [^16] https://novasky-ai.github.io/posts/sky-t1/. Nonetheless, by comparing our method with the SFT baseline, we observe clear improvements, which demonstrate the efficacy of our approach in enhancing the reasoning abilities of small models in various scenarios.
Dataset/Model Zero-shot SFT Ours
AIME2024 10.0 16.7 20.0
MATH-500 73.6 73.2 77.0
GPQA Diamond 33.3 28.8 36.9
LiveCodeBench V2 30.7 20.9 33.3
A.4 Results on Larger CoT Dataset
Dataset/Model Zero-shot SFT Ours
AIME2024 10.0 31.3 43.3
MATH-500 73.6 83.0 88.4
GPQA Diamond 33.3 42.4 42.9
LiveCodeBench V2 30.7 39.9 46.4
A.5 Design Choice of the Critic
An initial, straightforward approach is to employ $\pi_{\text{base}}$ as the Critic. However, owing to the small model’s limited reasoning capability, it consistently faces difficulties in distinguishing the difficulty levels of CoTs effectively within our datasets. Note that for CoTs rated as “easy” or “hard”, the CoT is either overly concise (omitting necessary steps) or excessively complex, rendering it unintelligible to the small model and preventing it from following the chain to arrive at the correct answer. Under these circumstances, it is clearly unreasonable to require the small model to classify the CoT difficulty that it cannot comprehend effectively.
Another intuitive CoT evaluation approach is to input the problem and its corresponding CoT into the small model and verify whether the model can arrive at the correct answer. However, applying this method directly would only partition CoT processes into “correct” or “incorrect” categories. For incorrect CoTs, this binary classification fails to distinguish the root cause of errors (i.e., whether the CoT is overly simplified or overly complex), which is critical for determining appropriate refinement strategies (e.g., expansion for overly simplified processes vs. simplification for overly complex ones).
Consequently, we utilize the larger and stronger LLM used in both the Rethinker and the Verifier (referred to as $\pi_{\text{large}}$) to act as the Critic. This involves guiding the large model to simulate the cognitive approach of the smaller model, $\pi_{\text{base}}$. The prompt template of the Critic is shown in Table 13. This setting is akin to educational practices, where professors, instead of students, customarily curate academic content across a spectrum of difficulty levels due to their broader knowledge base. As shown in Table 4, the experiments clearly demonstrate the superior evaluative proficiency of the large model, confirming its advantage in categorizing CoT complexity from the perspective of the smaller model efficiently.
Appendix B Case Studies
Case studies are presented in Tables 11 and 12.
Appendix C Prompt Templates
Prompt templates for the Critic, Rethinker, and Verifier in our CRV system are shown in Table 13.
@@ Appendix C Prompt Templates
Incorrect reasoning process $\text{Area}=\frac{1}{2}(\text{Sum of sides})=\frac{1}{2}(3+4+5)=6$
@@
Your task is to assess the given reasoning process from the perspective of a small language model (e.g.,
7B).
Specifically, determine whether the reasoning process provides sufficient detail for a small model to
solve the problem, or whether it is too terse (i.e., lacking critical details) or too complex (i.e., containing
@@
Complexity Definitions (from the perspective of a small model):
- Easy: The reasoning process is overly terse; it omits essential details that a small model needs to solve
the problem.
@@
- Hard: The reasoning process is overly complex, with extraneous or convoluted steps that could hinder
a small model to follow it.
@@
You must output exactly one word: easy, medium, or hard.
@@
(easy) Given a problem ,its correct answer and its terse reasoning process, your task is to extend the reasoning
process by adding necessary details and intermediate steps so that a small language model (e.g., a 7B
model) can follow the extended reasoning process to solve the problem.
You should add necessary steps and details based on the correct answer.
@@
(hard) Given a problem, its correct answer and its overly complex reasoning process, your task is to simplify the
reasoning process so that a small language model (e.g., a 7B model) can reliably follow the steps to solve
the problem.
You should remove redundancies or use simpler method on the basis of correct
answer.
@@
Your task is to assess a given reasoning process based on a problem and its correct answer.
Specifically, determine whether the reasoning process is sufficient and accurate for you to reach the correct
answer.
If the reasoning process correctly guides you to derive the the correct answer, output YES.
If the reasoning process fails to guide you to the correct answer, output NO.
You must output exactly one word: YES or NO.
Rethinker You are an assistant who is skilled at converting correct reasoning processes to incorrect reasoning
(incorrect) processes. Given a problem, its answer and its correct reasoning process, your task is to corrupt the correct
reasoning process by introducing logical fallacies and misleading steps, so that a small language model (e.g.,
a 7B model) cannot follow the incorrect reasoning process to solve the problem.
You must output ONLY the incorrect reasoning process with no additional explanation or commentary.
Table 13: Prompt templates for the CRV+CogPO framework.